Consider the following balanced equation:

3NH4NO3 + Na3PO4 → (NH4)3PO4 + 3NaNO3

30 g NH4NO3 and 50 g Na3PO4 are allowed to react.

-Which reactant is the limiting reactant?


1
Expert's answer
2021-11-30T18:50:01-0500

3NH4NO3 + Na3PO4 → (NH4)3PO4 + 3NaNO3

U

Mole ratio for NH4NO3 : Na3PO4

Is 3 : 1

Moles of NH4NO3= mass/Molecular mass

Molecular mass= 14+ 4 +14 +16×3= 80

Moles = 30÷80 = 0.375


Moles of Na3PO4 = mass ÷ molar mass

Molecular mass of Na3PO4 = 23 × 3 + 31 × 16 × 4= 164

Moles of Na3PO4 = 50÷164 = 0.305


The limiting reactant is NH4NO3